OpenStudy (anonymous):

evaluate the indefinite integral. sec(x)^21tan(x)dx

OpenStudy (anonymous):

alright , we have :\[\int\limits \sec^{21}(x)\tan(x) dx\] you can write it down as : \[\int\limits \sec^{20}(x) \tan(x)\sec(x) dx\] now let : u = sec(x) du = tan(x)sec(x) so now you'll get: \[\int\limits u^{20} du = \frac{u^{21}}{21} + c\] \[= \frac{\sec^{21}(x)}{21} + c\] correct me if I'm wrong ^_^